OpenStudy (anonymous):

the circumference of a circle is 10 (wierd symbol) inches find is radius and area. The radius of the circle is ( ) in.

OpenStudy (anonymous):

weird symbol?

OpenStudy (anonymous):

looks like a cursive r

OpenStudy (anonymous):

can you draw it?

OpenStudy (anonymous):

Do you mean pi? As in π?

OpenStudy (lgbasallote):

cursive r looks like \(\large \pi\)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i feel so stupid

OpenStudy (anonymous):

well, you should.

OpenStudy (anonymous):

new to algebra

OpenStudy (anonymous):

circumference = 2 pi r area = pi r^2 where r is the radius

OpenStudy (anonymous):

so if circumference = 10pi 10pi / 2pi = r correct?

OpenStudy (anonymous):

C= 2πr 10π = 2πr To find r, isolate it. \[\frac{10π}{2π} = \frac{2πr}{2π}\]\[\frac{5\cancel{10π}}{\cancel{2π}} = \frac{\cancel{2πr}}{\cancel{2π}}\]\[r = 5\]Now, the formula for area is \(πr^{2}\).Plug the radius into the formula and solve for the area. \[A = πr^{2}\]\[A = π(5)^{2}\]\[A = ?\]

OpenStudy (anonymous):

r is correct but what is the area

OpenStudy (anonymous):

I gave yo what you need. Now just follow the formula and solve for it.

OpenStudy (anonymous):

if you have r and area = pi r^2 what can you do?

OpenStudy (anonymous):

this is my first week in college algebra and i have no clue what i am doing

OpenStudy (anonymous):

college? you must be joking.

OpenStudy (anonymous):

This is barely algebra I and more like basic arithmetic. Do you know how to square something? \[A = πr^{2}\]\[A = π(5)^{2}\]

OpenStudy (anonymous):

r = 5 area = pi r^2 put r into the area formula

OpenStudy (anonymous):

times it by itself

OpenStudy (anonymous):

Yes. THat's what we're asking you to do.

OpenStudy (anonymous):

so 25

OpenStudy (anonymous):

lol and if i might ask, what college do you attend?

OpenStudy (anonymous):

Yes. A = 25π If you want an exact answer, this is it. If you want an approximate answer, it's A = 3.14(25).

OpenStudy (anonymous):

university of Phoenix